OpenStudy (anonymous):

What is the vertex of y = |x + 1|?

OpenStudy (cwrw238):



OpenStudy (cwrw238):

y = }x+ 1| is the part above the x-axis the vertex is where y = x+ 1 cuts the x axis ( when y = 0)

OpenStudy (cwrw238):

the part of y = x + 1 below x-axis is reflected in x-axis giving y = |x+1|